Matrix A is 2x2 matrix

Let A = [(1) (4) (3) (5)]

1.a Find A^-1 by using row operations
1.b Use question 1.a to Express A as a product of elementary matrices
1
Expert's answer
2021-08-08T16:52:56-0400

1. a


"A=\begin{pmatrix}
 1& 4\\
 3 & 5
\end{pmatrix}"

"\det A=\begin{vmatrix}
 1 & 4 \\
 3 & 5
\end{vmatrix}=5-12=-7\not=0=>A^{-1}\ exists"

"A^{-1}=\dfrac{1}{-7}\begin{pmatrix}
 5 & -4 \\
 -3 & 1
\end{pmatrix}"

"A^{-1}=\begin{pmatrix}
 -5/7& 4/7 \\
 3/7 & -1/7
\end{pmatrix}"

2. Augment the matrix with the identity matrix:


"\begin{pmatrix}
 1 & 4 & & 1 & 0 \\
 3 & 5 & & 0 & 1
\end{pmatrix}"

"R_2=R_2-3R_1"


"\begin{pmatrix}
 1 & 4 & & 1 & 0 \\
 0 & -7 & & -3 & 1
\end{pmatrix}"

"R_2=R_2\/(-7)"


"\begin{pmatrix}
 1 & 4 & & 1 & 0 \\
 0 & 1 & & 3/7 & -1/7
\end{pmatrix}"

"R_1=R_1-4R_2"


"\begin{pmatrix}
 1 & 0 & & -5/7 & 4/7 \\
 0 & 1 & & 3/7 & -1/7
\end{pmatrix}"

"A^{-1}=\begin{pmatrix}
 -5/7& 4/7 \\
 3/7 & -1/7
\end{pmatrix}"




"R_2=R_2+3R_1"


"\begin{pmatrix}
 1 & 0 \\
 3 & 1
\end{pmatrix}"

"R_2=-7R_2"


"\begin{pmatrix}
 1 & 0 \\
 0 & -7
\end{pmatrix}"

"R_1=R_1+4R_2"


"\begin{pmatrix}
 1 & 4 \\
 0 & 1
\end{pmatrix}"



"A=\begin{pmatrix}
 1& 4\\
 3 & 5
\end{pmatrix}=\begin{pmatrix}
 1 & 0 \\
 3 & 1
\end{pmatrix}\begin{pmatrix}
 1 & 0 \\
 0 & -7
\end{pmatrix}\begin{pmatrix}
 1 & 4 \\
 0 & 1
\end{pmatrix}"